OpenStudy (anonymous):

Sasha solved an equation. Her work is shown below. Explain Sasha's error. Then, show the correct way to solve. h divided by seven equals six. Add seven to both sides: h divided by seven plus seven equals six plus seven, so h equals thirteen.

OpenStudy (anonymous):

h/7 = 6 says the question. if we add 7 to both sides, then (h/7) + 7 = 6+7.. the error here: